OpenStudy (anonymous):

x^2-4x=1 what number must you add to complete this square?

OpenStudy (anonymous):

Please help it would be appreciated:)

OpenStudy (anonymous):

Remember,Half and square

OpenStudy (zarkon):

4 \[x^2-4x+4=(x-2)^2\]

OpenStudy (anonymous):

think "half of -4 is..." i get -2 then think "(-2) squared is..." 4. so you need to add 4

OpenStudy (anonymous):

Thank you and yes i remember imranmeah91.. but does that apply to all equations?

myininaya (myininaya):

\[x^2-4x+(\frac{-4}{2})^2=1+(\frac{-4}{2})^2\] \[x^2-4x+4=1+4\] \[(x-2)^2=5\] \[x-2=\pm \sqrt{5}\] now just add 2 to both sides

myininaya (myininaya):

\[x=2 \pm \sqrt{5}\]

OpenStudy (anonymous):

Yes, all, as long as there is 1 next to x^2 terms

OpenStudy (anonymous):

for example x^2 + 8x= -3? run me through that one please.. im sorry

OpenStudy (anonymous):

half of 8 is 4 square of 4 is 16 So 16

OpenStudy (anonymous):

okay thank you:) im a new fan

OpenStudy (anonymous):

x^2-4x = 1 x^2 - 4x -1 = 0 you want to have: (x-2)^2 = (2-x)^2 = x^2 + 4 - 4x so you must add 5 to -1 X^2 -4= 1+5